OpenStudy (anonymous):

The population P(t) of a species satisfies the logistic differential equation dP/dt=P(2-P/5000), where the initial population P(0)=3000 and t is the time in years. What is limit of P(t) as t approaches infinity?

OpenStudy (anonymous):

@dumbcow @ganeshie8 @iambatman

ganeshie8 (ganeshie8):

separate variables and solve the DE first

ganeshie8 (ganeshie8):

\[\dfrac{dP}{dt} = P\left(2-\dfrac{P}{5000}\right) \] \[ \int \dfrac{dP}{P\left(2-\dfrac{P}{5000}\right)} =\int dt \]

ganeshie8 (ganeshie8):

partial fractions on left side + solving explicitly gives u : \[P = \dfrac{10000 e^{2t}}{C + e^{2t}}\] take the limits as \(t \to \infty\)

OpenStudy (dumbcow):

separate variables \[\frac{dP}{P(10,000-P)} = \frac{dt}{5000}\] integrate \[\frac{1}{2}(\ln P - \ln (10,000-P)) = \frac{t}{5000} + C\] \[\frac{P}{10,000-P} = C e^{t/2500}\] solve for C using initial value C = 3/7 \[P = \frac{(30,000/7 e^{t/2500}}\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

for some reason there's this rule that it's always the denominator which here is 5000